### v1

# Fake News in Sheep’s Clothing: Robust Fake News Detection Against LLM-Empowered Style Attacks

Abstract.
It is commonly perceived that online fake news and reliable news exhibit stark differences in writing styles, such as the use of sensationalist versus objective language. However, we emphasize that style-related features can also be exploited for style-based attacks. Notably, the rise of powerful Large Language Models (LLMs) has enabled malicious users to mimic the style of trustworthy news outlets at minimal cost. Our analysis reveals that LLM-camouflaged fake news content leads to substantial performance degradation of state-of-the-art text-based detectors (up to 38% decrease in F1 Score), posing a significant challenge for automated detection in online ecosystems. To address this, we introduce SheepDog, a style-agnostic fake news detector robust to news writing styles. SheepDog achieves this adaptability through LLM-empowered news reframing, which customizes each article to match different writing styles using style-oriented reframing prompts. By employing style-agnostic training, SheepDog enhances its resilience to stylistic variations by maximizing prediction consistency across these diverse reframings. Furthermore, SheepDog extracts content-focused veracity attributions from LLMs, where the news content is evaluated against a set of fact-checking rationales. These attributions provide supplementary information and potential interpretability that assist veracity prediction. On three benchmark datasets, empirical results show that SheepDog consistently yields significant improvements over competitive baselines and enhances robustness against LLM-empowered style attacks.
Fake News; Large Language Models; Adversarial Robustness
ccs: Information systems Data miningccs: Computing methodologies Natural language processing
1. Introduction
Figure 1. A motivating example of LLM-empowered style attacks on text-based fake news detectors, where fake news is camouflaged with the style of reliable news publishers.
Automated fake news detectors are central to safeguarding integrity and harmony in the digital age (Nguyen et al., 2020; Shu et al., 2019; Wu and Hooi, 2023; Zhou et al., 2020). Style-related signals embedded in news articles serve as a crucial indicator of news veracity. For instance, reputable news sources uphold journalistic integrity, emphasize accuracy and fact-checking, and maintain a balanced tone (Bachmann et al., 2021). In stark contrast, unreliable outlets often employ sensationalism, lack credible sources, and may exhibit partisan biases (Higdon, 2020). Building upon this, prior advances have extracted sentiment features (Ajao et al., 2019; Zhang et al., 2021) to complement and enhance the detector, and highlighted the significance of writing style between hyperpartisan news and well-balanced mainstream reporting (Potthast et al., 2018).
While style-related features are invaluable for discerning news veracity, they also offer a direct avenue for malicious users to conduct style-based attacks. This problem is exacerbated by the advent of powerful Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022), whose unprecedented capabilities for reasoning and generative tasks (Wei et al., 2022; Zheng and Zhan, 2023) reduces the gap between machine-generated and human-written news. Consequently, malicious actors now possess the capability to mimic the style of reputable news sources, in an attempt to evade automated detection. As shown in Figure 1, using a style-oriented prompt (i.e., “style of The New York Times”), LLM-camouflaged fake news successfully bypasses a RoBERTa (Liu et al., 2019)-based fake news detector.
The impact of style-based attacks on fake news detectors has received limited attention, despite prior investigations into detector vulnerability related to attacks regarding social engagement (Chen and Shu, 2023), word-level perturbations (Koenders et al., 2021), and machine-generated malicious comments (Le et al., 2020). To assess the robustness of text-based fake news detectors, we introduce a series of style-based adversarial attacks, specifically by tailoring news articles to adversarial writing styles (details provided in Section 4.1). Our experiments, as summarized in Table 1, reveal a significant performance degradation in state-of-the-art text-based detectors, with some experiencing up to a 38% decline in F1 Score. These detectors, well-adapted to real news from trustworthy sources and fake news from unreliable sources, struggle to address the emerging real-world challenge where news content is presented in various styles.
Motivated by the style-related vulnerability observed in state-of-the-art text-based detectors, we present SheepDog — a robust, style-agnostic fake news detector that effectively identifies deceptive content, even when concealed within the LLM-empowered ”sheep’s clothing”. Built upon a pretrained language model (LM) backbone, which can be fine-tuned to capture task-specific salient features, and leveraging the strong zero-shot reasoning and generative capabilities of LLMs, SheepDog capitalizes on the strengths of both (LLM: versatile; fine-tuned LM: specialized).
Our key ideas for overcoming style attacks are: (1) style-agnostic training, in which predictions across style-diverse news reframings are regulated to ensure maximal consistency, and (2) content-focused veracity attributions, in which textual insights on news veracity are extracted from an LLM to inform veracity predictions.
The style-agnostic nature of SheepDog stems from its utilization of LLM-empowered news reframings. To encompass a wide spectrum of real-world news writing styles, we expand our training samples with style-diverse expressions, and ensure coherence across different expressions to establish style invariance. Specifically, through style-oriented prompts, we exploit the impressive text generation capabilities of LLMs to generate reframings for each news article, introducing a diversity of styles. Then, we maximize the consistency in the model’s veracity predictions across the original article and its various reframings. This strategy assists SheepDog to discount style-related features, which enables it to capture style-agnostic signals from the news content.
To train a content-focused detector that is robust against stylistic variations, we further propose to distill the veracity-related knowledge from LLMs, specifically via eliciting content-focused attributions from inference APIs. We instruct the LLM to assess each article based on a defined set of content-focused fact-checking rationales (for comprehensive details, refer to Section 5.3). The resulting textual attributions serve to identify indicators of fake news. By converting these attributions into precise pseudo-labels, we introduce additional weak supervision that steers SheepDog towards learning a robust, style-agnostic representation of news. Our approach leverages these attribution-level predictions not only to facilitate style-agnostic training but also to potentially offer explainability into the veracity of news articles.
Overall, we make the following three contributions:
  • [leftmargin=*]
  • Empirical Finding: We present a novel finding on the style-related vulnerability of state-of-the-art text-based fake news detectors to LLM-empowered style attacks.
  • LLM-Empowered Style Robustness: We introduce SheepDog, a style-agnostic fake news detector that achieves robustness through a combination of style-agnostic training and content-focused veracity attributions.
  • Effectiveness: Through extensive experiments, we demonstrate that leveraging LLM-empowered diverse news reframings significantly enhances the performance of representative language models (LMs) across three benchmark datasets.
2. Related Work
Fake News Detection. Automated fake news detection has been widely explored using a wide range of neural architectures (Devlin et al., 2019; Liu et al., 2019; Pelrine et al., 2021; Shu et al., 2019; Zhou et al., 2020). Apart from extracting lexical (Rashkin et al., 2017) and sentiment features (Potthast et al., 2018) within the news article text, many methods incorporate auxiliary features to supplement veracity prediction, including user comments (Shu et al., 2019), news environments (Sheng et al., 2022), knowledge bases (Cui et al., 2020; Dun et al., 2021), temporal patterns from users (Ruchansky et al., 2017), and social graphs (Lu and Li, 2020; Nguyen et al., 2020; Wu and Hooi, 2023). Recent studies also seek to address challenges including temporal shift (Hu et al., 2023), entity bias (Zhu et al., 2022a) and domain shift (Nan et al., 2021, 2022; Zhu et al., 2022b) in fake news detection scenarios. In this work, we adopt a text-based perspective and focus on news article texts, specifically on the stylistic variations in the news article under LLM-empowered style attacks.
Adversarial Attack on Fake News Detectors. Investigating the vulnerabilities of fake news detectors is central to improving their real-world applicability. Hence, existing efforts (He et al., 2021a; Horne et al., 2019; Koenders et al., 2021; Lyu et al., 2023; Wang et al., 2023; Zhou et al., 2019) have studied the impact of different attacks from multiple aspects, including manipulation of social engagements (Lyu et al., 2023; Wang et al., 2023) and user behavior (He et al., 2021a), fact distortion (Koenders et al., 2021), subject-object exchange (Zhou et al., 2019), and blocking of data availability (Horne et al., 2019). However, the impact of writing styles remains underexplored. To bridge this gap, we investigate the robustness of text-based detectors against LLM-empowered style attacks, and propose a style-agnostic solution.
Large Language Models (LLMs). LLMs (OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023), with billions of parameters, have showcased remarkable reasoning capabilities, even matching or surpassing human performance in some scenarios (Wei et al., 2022; Zheng and Zhan, 2023). Benefiting from this success, recent work extracts knowledge from LLMs to facilitate a wide range of tasks, such as visual classification (Menon and Vondrick, 2023) and node classification (He et al., 2023). On a related front, LLM-generated text detection (Henrique et al., 2023; Muñoz-Ortiz et al., 2023) has also attracted increasing attention. Recent investigations further explore LLM-generated misinformation (Epstein et al., 2023; Huang et al., 2023; Zellers et al., 2019; Chen and Shu, 2023), and leverage LLMs for program-guided fact-checking (Pan et al., 2023). Although we also prompt an LLM to generate adversarial articles, we aim for style-agnostic detection that yields robust predictions across style-diverse news expressions. Moreover, existing work on LLM-generated misinformation mainly focuses on analysis of LLM-induced impacts and news article generation, which is orthogonal to our contributions. In this work, we focus on improving the detector’s style robustness by eliciting news reframings and content-focused veracity attributions from LLMs.
3. Problem Definition
Let $\mathcal{D}$ be a news dataset consisting of $N$ questionable news pieces, denoted as $p_{1},p_{2},\dots,p_{N}$. Among the news pieces, $\mathcal{P}_{L}\subset\mathcal{D}$ is a set of labeled news articles. Each news article in $\mathcal{P}_{L}$ is assigned a ground-truth veracity label $y$. In line with prior work (Shu et al., 2019; Zhang et al., 2021; Zhou et al., 2020), $y$ is a binary label that represents either real news or fake news.
As we focus on style-related issues, we consider a text-based setting. Formally, the problem can be defined as follows:
Problem 1 (Text-Based Fake News Detection).
Given a news dataset $\mathcal{D}$ with training labels $\mathcal{Y}_{L}$, the goal is to predict the veracity labels of unlabeled news pieces $\mathcal{P}_{U}=\mathcal{D}\setminus\mathcal{P}_{L}$.
4. LLM-Empowered Style Attacks
In this section, we establish a series of LLM-empowered style attacks, and conduct preliminary analysis to assess the robustness of state-of-the-art text-based fake news detectors.
4.1. Attack Formulation
The impressive capabilities of LLMs (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022) have enabled malicious users to disguise fake news with restyling prompts, resulting in camouflaged articles that closely resemble reliable sources. Here, we explore a direct form of style-based attack utilizing news publisher names (e.g., ”The New York Times”). These names possess distinct styles that can be readily adopted by producers of fake news, making them a likely occurrence in real-world scenarios.
To replicate adversarial situations where news pieces are restyled in relation to various publishers, we manipulate the styles of both trustworthy and unreliable news. Specifically, in our test data, we utlize an LLM to rephrase real news in the style of tabloids, and fake news in the style of mainstream sources. Our general prompt format is shown as follows: {tcolorbox}[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article using the style of [publisher name]: [news article] Based on publisher popularity, in the place of [publisher name], we select “National Enquirer” to camouflage real news, and “The New York Times” to camouflage fake news. These LLM-camouflaged test articles are then employed to evaluate the resilience of a detector against style-based attacks, as illustrated in Figure 1.
4.2. Style-Related Detector Vulnerability
Automated fake news detection becomes increasingly difficult with LLM-empowered style attacks. In this subsection, we conduct preliminary analysis on real-world news to evaluate the influence of writing styles on text-based detectors. Our analysis is based on the FakeNewsNet (Shu et al., 2020) benchmark (consisting of PolitiFact and GossipCop datasets) and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). Further details regarding these benchmark datasets can be found in Table 8 and Section 6.1.1. Specifically, we seek to address the following question: To what extent can text-based fake news detectors withstand LLM-empowered style attacks?
In Table 1, we examine $13$ representative text-based detectors under both original and adversarial settings (detailed method descriptions are provided in Section 6.1.2). These detectors employ diverse model architectures, including Convolutional Neural Networks (CNNs) (Zhou et al., 2020), Recurrent Neural Networks (RNNs), (Shu et al., 2019), Graph Neural Networks (GNNs) applied to document graphs (Vaibhav et al., 2019), Transformer-based LMs (Devlin et al., 2019; He et al., 2021b; Liu et al., 2019; Zhang et al., 2021; Schick and Schütze, 2021; Hu et al., 2022a; Xie et al., 2020), and LLMs (OpenAI, 2022; Ouyang et al., 2022; Touvron et al., 2023). We evaluate model robustness against LLM-empowered style attack based on performance on the adversarial test set formulated in Section 4.1. Our empirical results yield the following insights:
Observation 1 (Style-Related Vulnerability of Fake News Detectors).
State-of-the-art text-based fake news detectors are susceptible to LLM-empowered style attacks. This susceptibility results in substantial performance degradation,with an F1 Score decline of up to 38.3% on the adversarial test set.
Observation 2 (Insufficiency of LLMs as Zero-Shot Fake News Detectors).
LLMs, despite their impressive zero-shot capabilities as general-purpose foundation models, exhibit inferior detection performance compared to text-based fake news detectors and pre-trained LMs fine-tuned specifically for fake news detection.
Our two findings reveal a critical challenge in achieving robustness again stylistic variations. Ideally, an effective fake news detector should prioritize content assessment over style. While the news content typically remains stable, style can undergo rapid changes in the dynamic online environment. Our observations, however, reveals a stark deviation from this ideal. Detectors overly influenced by style may struggle to reliably differentiate between real and fake news, and even powerful LLMs may prove inadequate for the specific demands of fake news detection. The accessibility for malicious users to manipulate style using LLMs further exacerbates this issue, emphasizing the pressing need for a style-agnostic approach.
Table 1. Under LLM-empowered style attacks, existing text-based fake news detectors suffer severe performance deterioration in terms of F1 Score (%). Performance degradation is computed on adversarial test set (O: original; A: adversarial).
Method PolitiFact GossipCop LUN
O A ($\downarrow$) O A ($\downarrow$) O A ($\downarrow$)
[…]
SentGCN (Vaibhav et al., 2019) 80.77 13.82 69.29 5.59 79.66 16.65
BERT (Devlin et al., 2019) 84.99 12.68 74.50 5.52 80.96 24.61
RoBERTa (Liu et al., 2019) 87.40 11.23 74.05 3.05 82.12 29.65
DeBERTa (He et al., 2021b) 86.30 11.73 73.80 2.85 83.67 30.34
UDA (Xie et al., 2020) 87.74 10.14 74.22 4.54 82.94 20.71
DualEmo (Zhang et al., 2021) 87.76 15.34 75.36 5.89 81.52 24.97
PET (Schick and Schütze, 2021) 85.51 11.02 74.63 3.08 83.66 31.08
KPT (Hu et al., 2022a) 87.70 13.26 74.23 2.63 84.06 31.83
GPT3.5 (OpenAI, 2022) 69.61 27.48 56.30 16.71 79.97 20.34
InstructGPT (Ouyang et al., 2022) 64.59 20.69 50.38 9.13 68.16 11.39
Llama2-13B (Touvron et al., 2023) 63.15 29.91 53.54 27.75 70.97 38.33
Figure 2. Overview of our SheepDog style-agnostic fake news detector.
5. Proposed Approach
Building upon our empirical findings regarding the style-related vulnerability of fake news detectors, and the limitations of LLMs as zero-shot detectors, we present SheepDog, a style-agnostic detector that reliably assesses news veracity through the utilization of LLMs. As illustrated in Figure 2, the robustness of SheepDog is obtained from two core objectives in a multi-task learning paradigm: (1) style-agnostic training, which flexibly adapts a pre-trained LM to the fake news classification objective, and ensures consistent veracity predictions across a diverse array of news reframings; and (2) content-focused veracity attributions, which extracts veracity-related insights from LLMs to inform model predictions.
To enhance usability and composability, we design our method to be simple and modular, allowing it to be used with any language model backbone.
5.1. LLM-Empowered News Reframing
In light of our Observation 1, it becomes evident that text-based fake news detectors, when exclusively trained on style-consistent news articles, may exhibit limited real-world adaptability. To overcome this limitation, our key idea is to diversify the training data through a process we refer to as reframing, which presents each news articles in various styles.
Our reframing strategy is driven by two sub-goals: (1) ensuring coherence in generated texts, and (2) covering a diverse range of styles. LLMs inherently exhibit both desirable properties. Hence, for each training article, we generate a series of prompts, each consisting of the news article and a style-oriented reframing instruction. The general prompt format is represented as follows, with a detailed example provided in Table 9:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a / an [specified] tone: [news article] To generate news expressions that simulate both reliable and unreliable sources, we establish a set of four general style-oriented adjectives for the prompt: ”objective and professional” and ”neutral” to emulate reliable sources, and ”emotionally triggering” and ”sensational” for unreliable ones.
During the training phase, for labeled news article $p\in\mathcal{P}_{L}$, we randomly select one reliable-style reframing prompt and one unreliable-style reframing prompt to generate diverse expressions. Through querying the LLM, we obtain two corresponding reframings: one reliable-style reframing denoted as $p_{R}$, and one unreliable-style reframing denoted as $p_{F}$.
5.2. Style-Agnostic Training
A robust fake news detector should possess the capability to discern the veracity of news articles based on their content, without being influenced by their stylistic attributes. To this end, we introduce a style alignment objective that ensures close alignment among the veracity predictions of news article $p$, its reliable-style reframing $p_{R}$, and its unreliable-style reframing $p_{F}$. This objective is derived as follows.
Let $\mathcal{M}$ be a pre-trained Language Model (LM) such as RoBERTa (Liu et al., 2019). Employing an LM as the backbone of the detector offers advantages, as LMs can be readily fine-tuned for the fake news detection task, which enables them to effectively extract salient task-specific features from the news content. Through $\mathcal{M}$, based on $p$, $p_{R}$ and $p_{F}$, we acquire article representations $\mathbf{h}\in\mathbb{R}^{d}$, and reframing representations $\mathbf{h}_{R},\mathbf{h}_{F}\in\mathbb{R}^{d}$:
(1) $\mathbf{h}_{p}=\mathcal{M}(p),\quad\mathbf{h}_{R}=\mathcal{M}(p_{R}),\quad\mathbf{h}_{F}=\mathcal{M}(p_{F}).$
Subsequently, we apply a Multi-Layer Perceptron (MLP) to these representations, and obtain veracity predictions $\tilde{\mathbf{y}},\tilde{\mathbf{y}}_{R},\tilde{\mathbf{y}}_{F}\in\mathbb{R}^{2}$ for $p$, $p_{R}$ and $p_{F}$:
(2) $\tilde{\mathbf{y}}=\text{MLP}_{pred}(\mathbf{h}),\quad\tilde{\mathbf{y}}_{R}=\text{MLP}_{pred}(\mathbf{h}_{R}),\quad\tilde{\mathbf{y}}_{F}=\text{MLP}_{pred}(\mathbf{h}_{F}).$
Each of $\tilde{\mathbf{y}},\tilde{\mathbf{y}}_{R},\tilde{\mathbf{y}}_{F}$ contain two logits that correspond to the real and fake classes, respectively.
Despite differences in style, the fundamental news content remains consistent across the original news article $p$ and its reframings $p_{R}$ and $p_{F}$. Ideally, $p_{R}$ and $p_{F}$ should yield the same predicted label as $p$. To facilitate this, we formulate the following style alignment loss defined as follows:
(3) $\mathcal{L}_{\text{style}}=\textsf{MEAN}\left(\mathcal{L}_{1}(\tilde{\mathbf{y}}_{R},\tilde{\mathbf{y}}),\;\mathcal{L}_{1}(\tilde{\mathbf{y}}_{F},\tilde{\mathbf{y}})\right),$
where $\mathcal{L}_{1}$ represents the The Kullback-Leibler (KL) divergence loss.
While aligning the predictions of $p_{R}$ and $p_{F}$ with $p$, it is crucial to ensure accurate predictions for $p$. Therefore, we also incorporate a fake news detection loss:
(4) $\mathcal{L}_{\text{news}}=\mathcal{L}_{2}(\tilde{\mathbf{y}},y),$
where $\mathcal{L}_{2}$ represents the standard cross entropy (CE) loss.
5.3. Content-Focused Veracity Attributions
In addition to tuning the LM backbone with the style alignment objective for robust predictions, to capture the veracity-related signals directly linked to the news content, we further propose to integrate external knowledge and reasoning to inform veracity predictions. General-purpose LLMs, known for their remarkable zero-shot reasoning capabilities (Huang and Chang, 2023), provides a promising solution.
Specifically, we generate content-focused veracity attributions by prompting an LLM. Our prompt consists of the news article text and a predefined set $\mathcal{C}$ of content-oriented fact-checking rationales (e.g., “lack of credible sources” and “false or misleading information”; we present the detailed prompt template and an example in Appendix B.3 and Table 10). Through this prompt, we instruct the LLM to provide textual attributions indicative of news veracity (i.e., identifying characteristics associated with fake news). This process efficiently extracts the LLM’s reasoning capabilities and general knowledge, presenting it in a text format. The resulting textual attributions, in turn, enriches our model with additional veracity-informed knowledge. The general prompt format is shown as follows:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Article: [news article]
Question: [given a list of fact-checking rationales, ask the LLM to identify content-related characteristics associated with fake news] Upon querying the LLM, we obtain a list of textual veracity attributions based on the input article. These attributions are then transformed into $|\mathcal{C}|$-dimensional pseudo-labels, where the attribution to each rationale in $\mathcal{C}$ is represented by a distinct binary label. For a given news article $p\in\mathcal{P}_{L}$, this process yields pseudo-labels $\mathbf{s}\in\mathbb{R}^{|\mathcal{C}|}$, which provides supplementary veracity-related information. Similarly, for reframings $p_{R}$ and $p_{F}$, we obtain pseudo-labels $\mathbf{s}_{R},\mathbf{s}_{F}\in\mathbb{R}^{|\mathcal{C}|}$, respectively. Notably, as $\mathcal{C}$ is focused on indicators of fake news, the pseudo-labels for real news and its reframings are uniformly set to all-zeros.
To incorporate the veracity-informed knowledge from these content-focused veracity attributions into our detector, we introduce a multi-label attribution prediction objective. As shown in Eq. 1, we learn news representations $\mathbf{h},\mathbf{h}_{R},\mathbf{h}_{F}\in\mathbb{R}^{d}$ for news article $p$ and its reframings $p_{R}$ and $p_{F}$, respectively, using a pre-trained LM $\mathcal{M}$. Then, the attribution-level prediction scores $\tilde{\mathbf{s}},\tilde{\mathbf{s}}_{R},\tilde{\mathbf{s}}_{F}\in\mathbb{R}^{|\mathcal{C}|}$ are computed through another MLP:
(5) $\tilde{\mathbf{s}}=\text{MLP}_{attr}(\mathbf{h}),\quad\tilde{\mathbf{s}}_{R}=\text{MLP}_{attr}(\mathbf{h}_{R}),\quad\tilde{\mathbf{s}}_{F}=\text{MLP}_{attr}(\mathbf{h}_{F}).$
The veracity attribution loss is then defined as:
(6) $\mathcal{L}_{\text{attr}}=\textsf{MEAN}\left(\mathcal{L}_{3}(\hat{\mathbf{s}},\;\mathbf{s}),\mathcal{L}_{3}(\hat{\mathbf{s}_{R}},\;\mathbf{s}_{R}),\mathcal{L}_{3}(\hat{\mathbf{s}_{F}},\;\mathbf{s}_{F})\right),$
where $\mathcal{L}_{3}$ represents the binary cross entropy (BCE) loss, and $\hat{\mathbf{s}}$ denotes the sigmoid-transformed scores in $\tilde{\mathbf{s}}$ corresponding to each rationale.
5.4. Final Objective Function of SheepDog
By enforcing consistency among style-diverse news reframings and exploiting the content-focused attributions from the LLM, the final objective function of SheepDog is defined as a linear combination of the the style alignment loss (Eq. 3), the news classification loss (Eq. 4), and the veracity attribution loss (Eq. 6):
(7) $\mathcal{L}=\mathcal{L}_{\text{style}}+\mathcal{L}_{\text{news}}+\mathcal{L}_{\text{attr}}.$
Our proposed SheepDog detector is an end-to-end framework where we simultaneously learn the style-agnostic news veracity predictor and the content-focused attribution predictor.
[…]
Method PolitiFact GossipCop LUN
A B C D A B C D A B C D
[…]
SentGCN 66.95 62.50 69.54 65.08 63.70 63.07 63.61 63.01 63.01 62.50 76.11 75.56
DualEmo 72.42 71.23 77.07 75.80 69.47 68.50 71.69 70.71 56.55 54.78 68.53 66.80
G2 BERT 72.31 71.37 77.23 76.24 68.98 68.17 71.95 71.11 56.35 54.61 68.50 66.74
RoBERTa 76.17 74.95 78.28 77.05 71.00 70.47 72.56 72.02 52.47 53.62 68.31 69.46
DeBERTa 74.57 74.36 80.60 80.35 70.95 71.15 72.51 72.71 53.33 55.45 67.16 69.27
UDA 77.60 75.57 79.21 77.17 69.68 69.33 72.16 71.80 62.23 61.80 68.25 67.80
PET 74.49 70.75 75.49 71.76 71.55 70.85 73.74 73.02 52.58 53.30 63.71 64.33
KPT 74.44 73.32 77.73 76.60 71.60 71.01 73.69 73.10 52.23 53.62 65.71 67.15
G3 GPT3.5 42.13 43.44 56.61 58.17 39.59 38.67 48.44 47.38 59.63 61.24 65.74 67.43
InstructGPT 43.90 43.90 54.21 54.21 41.25 40.18 44.26 43.12 56.77 57.15 58.93 59.32
Llama2-13B 33.24 34.48 53.64 55.45 25.79 26.06 37.07 37.40 32.64 33.00 50.81 51.33
Ours SheepDog $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ 81.24 $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
6. Experiments
In this section, we empirically evaluate SheepDog to investigate the following six research questions:
  • Robustness Against Style Attacks (Section 6.2): How robust is SheepDog against LLM-empowered style attacks?
  • Effectiveness on Unperturbed Articles (Section 6.3): How proficiently can SheepDog identify fake news within the original unperturbed test articles?
  • Adaptability to Different LM Backbones (Section 6.4): How does SheepDog’s performance vary with different representative LM backbones?
  • Ablation Study (Section 6.5): What are the respective impacts of style-agnostic training and content-focused attributions on SheepDog’s overall performance?
  • Stability Across Reframing Prompts (Section 6.6): Does SheepDog yield consistent improvements across diverse settings of news reframing prompts?
  • Case Study (Section 6.7): How can we interpret SheepDog’s rationale for fake news detection through its predictions on content-focused attributions?
6.1. Experimental Setup
6.1.1. Datasets
We evaluate our approach on three widely-used real-world benchmark datasets: the FakeNewsNet public benchmark (Shu et al., 2020), which consists of the PolitiFact and GossipCop datasets, and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). PolitiFact and LUN center on political discourse, while GossipCop focuses on celebrity gossip. PolitiFact and LUN center on political discourse, while GossipCop focuses on celebrity gossip.
The dataset description and statistics are delegated to Appendix A and Table 8. We randomly select $80$% of the articles as the training data, and the other $20$% as test data.
6.1.2. Baselines
We benchmark SheepDog against thirteen representative baseline methods, which can be categorized as:
[…]
Fine-tuned LMs (G2) adapts pre-trained LMs to the fake news detection task, and has proven effective in handling misinformation scenarios (Pelrine et al., 2021). In addition to three widely-recognized LMs, namely BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and DeBERTa (He et al., 2021b), we include UDA (Xie et al., 2020), a representative BERT-based model that employs diverse text augmentations to yield consistent model predictions against input noise. We also select two methods under the popular prompting paradigm: PET (Schick and Schütze, 2021), which converts textual inputs into cloze questions that contain a task description; and KPT, (Hu et al., 2022b) which expands the label word space with varied class-related tokens. For a fair comparison, as our proposed approach does not involve unlabeled articles, we implement UDA using consistency training on the supervised training data, and exclude the self-training and PLM ensemble components for PET. All methods in this category are implemented with base version LMs, in line with our approach.
LLMs (G3) generate textual veracity predictions in a zero-shot manner. We select three representative LLMs as baselines: GPT3.5 (OpenAI, 2022) (version name: gpt-3.5-turbo), InstructGPT (Ouyang et al., 2022) (version name: gpt-3.5-turbo-instruct), and Llama2-13B (Touvron et al., 2023) (version name: Llama-2-13b-chat-hf), with details delegated to Appendix B.1.
6.1.3. Implementation Details
We implement SheepDog and its variants based on PyTorch 1.10.0 with CUDA 11.1. We utilize pretrained RoBERTa-base weights from HuggingFace Transformers 4.13.0 (Wolf et al., 2020). The LM backbone for SheepDog was configured with a maximum sequence length of 512, a batch size of 4, and a learning rate of $2\times 10^{-5}$. We prompt GPT3.5 (version name: gpt-3.5-turbo) to generate news reframings, content-focused veracity attributions, and adversarial test articles. Detailed prompting descriptions and examples are delegated to Appendix B.2 and B.3. For SheepDog’s attribution prediction and veracity prediction, we employ two MLPs, each with a single layer (we also implement a variant with 2-layer MLPs in Section 6.5). Our model is fine-tuned for $5$ epochs. For the implementation of baseline methods, we adhere to the architectures and hyperparameters recommended by their respective authors.
We evaluate model performance using Accuracy (%) and macro-F1 Score (%). For all experiments except those involving LLMs, we report averaged metrics over 10 runs of each method to provide a comprehensive evaluation. In the case of LLM zero-shot predictions, we employ greedy decoding and conduct each experiment once.
6.2. Robustness Against Style Attacks
We establish a series of LLM-empowered style attacks to assess SheepDog’s robustness. Following our prompt template formulated in Section 4.1, in the place of [publisher name], we select “National Enquirer” and “The Sun” to camouflage real news, and “CNN” and “The New York Times” for fake news, according to publisher popularity. This yields $2\times 2=4$ distinct adversarial test sets, labeled as A through D in Table 3. Note that we report the results of SheepDog on adversarial set A in all other subsections.
Table 3. Notations and setup for the four style-based adversarial test sets in Section 6.2, denoted as A through D.
[publisher name] CNN The New York Times
National Enquirer A B
The Sun C D
[…]
6.3. Effectiveness on Unperturbed Articles
A desirable fake news detector should achieve both robustness and effectiveness. Our empirical results, detailed in Table 4, clearly demonstrate that SheepDog excels in this regard. When tested on the original, unaltered articles, SheepDog consistently matches (on PolitiFact and GossipCop) or surpasses (on LUN) the performance of the most competitive baseline, in terms of accuracy and F1 score. This showcases the framework’s proficiency in accurately discerning unreliable news from reliable ones, and implies SheepDog’s practical utility in real-world scenarios.
6.4. Adaptability to Different LM Backbones
To assess the flexibility of SheepDog, we evaluate the performance of SheepDog combined with three representative LM backbones: RoBERTa, BERT and DeBERTa. As demonstrated in Table 5, SheepDog substantially enhances the performance of the corresponding LM backbones. This adaptability highlights SheepDog’s robustness from style-agnostic training and content-focused attributions, showcasing its potential for flexible deployment in diverse contexts where different LMs may be preferred or more readily available. This finding further reinforces the benefits of SheepDog’s design with a LM backbone, which flexibly adapts it to task-specific needs.
[…]
Method PolitiFact GossipCop LUN
Acc. F1 Acc. F1 Acc. F1
[…]
SentGCN 81.11 80.77 69.38 69.29 80.07 79.66
DualEmo 87.78 87.76 75.51 75.36 81.78 81.52
BERT 85.22 84.99 74.60 74.50 81.13 80.96
RoBERTa 88.00 87.40 74.14 74.05 82.53 82.12
DeBERTa 86.33 86.30 73.86 73.80 84.01 83.67
UDA 87.77 87.74 74.28 74.22 83.02 82.94
PET 85.56 85.51 74.75 74.63 84.00 83.66
KPT 87.78 87.70 74.38 74.23 84.40 84.06
GPT3.5 71.11 69.61 61.49 56.30 80.67 79.97
InstructGPT 67.78 64.59 58.33 50.38 70.87 68.16
Llama2-13B 65.56 63.15 55.74 53.54 72.47 70.97
SheepDog 88.44 88.39 75.77 75.75 $*{}^{*}$ $*{}^{*}$
[…]
Method PolitiFact GossipCop LUN
RoBERTa 76.17 71.00 52.47
SheepDog-RoBERTa $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
BERT 72.31 68.98 56.35
SheepDog-BERT $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
DeBERTa 74.57 70.95 53.33
SheepDog-DeBERTa $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
6.5. Ablation Study
To gain deeper insights into the functioning of SheepDog and the role of its different components, we conduct an ablation study to evaluate the respective contributions of LLM-empowered news reframings and content-focused veracity attributions. We compare SheepDog with the following three model variants:
  • SheepDog w/ 2-layer MLP, which employs 2-layer MLPs with hidden size of $64$ as attribution detector and veracity detector.
  • SheepDog-R, which excludes style-diverse news reframings and the style-agnostic training component.
  • SheepDog-A, which omits content-focused veracity attributions and the attribution prediction component.
Table 6 shows the performance of different SheepDog variants on adversarial test sets C and D. The results suggest that: (1) SheepDog-R without news reframings only yields slight improvements compared with a fine-tuned RoBERTa model, which suggests the pivotal role of diverse reframings in SheepDog’s robustness. (2) Removing content-focused veracity attributions (SheepDog-A) results in inferior performance than SheepDog, which implies the effectiveness of extracting content-related knowledge from LLMs. (3) SheepDog, utilizing one single layer for veracity prediction and attribute predictions, slightly outperforms the variant employing 2-layer MLPs. This suggests that the expressiveness of LMs, harnessed through our proposed objectives, yields article representations that contains rich indicators for both attributions and veracity.
Table 6. Ablation of SheepDog demonstrates benefits of LLM-empowered news reframing (denoted as R) and content-focused veracity attributions (denoted as A) in F1 Score (%).
Method PolitiFact GossipCop LUN
SheepDog 80.99 74.45 85.63
  w/ 2-layer MLP 79.83 74.03 84.75
  - R 76.71 70.98 53.27
  - A 80.73 73.74 84.83
  RoBERTa 76.17 71.00 52.47
6.6. Stability Across Reframing Prompts
As detailed in Section 5.1, we utilize an LLM to generate two types of reframings for a given training article $p$: a reliable-style reframing denoted as $p_{R}$, and an unreliable-style reframing denoted as $p_{F}$. To assess SheepDog’s stability across different reframing prompts, we systematically examine its performance using four diverse combinations of prompts, denoted as R1 through R4.
Recall that we adopt the following template for news reframing:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a / an [specified] tone: [$p$]
For R1 through R4, the specified tones are defined as:
  • R1: $p_{R}$ “objective and professional”; $p_{F}$ “emotionally triggering”.
  • R2: $p_{R}$ “objective and professional”; $p_{F}$ “sensational”.
  • R3: $p_{R}$ “neutral”; $p_{F}$ “emotionally triggering”.
  • R4: $p_{R}$ “neutral”; $p_{F}$ “sensational”.
As shown in Table 7, SheepDog consistently demonstrates stable and significant improvements over the most competitive baseline. This validates the generalizability of our approach, suggesting its potential in effectively combating deceptive online information in the ever-evolving digital landscape. Notably, our SheepDog approach conducts sampling between two reliable-style reframings and two unreliable-style reframings, leading to better versatility.
Table 7. Across different sets of news reframing prompts, SheepDog demonstrates stable and significant improvements over the most competitive baseline (in F1 %). (“Baseline”: performance of the most competitive baseline.
Method PolitiFact GossipCop LUN
Baseline 77.60 71.60 66.34
SheepDog 80.99 74.45 85.63
  w/ R1 79.02 74.22 77.42
  w/ R2 79.93 74.16 86.18
  w/ R3 80.36 73.55 76.77
  w/ R4 79.71 74.01 85.55
Figure 3. SheepDog yields robust predictions against LLM-camouflaged fake news. Additionally, the top predicted veracity attribution points toward the false and misleading information in the news content, providing interpretability.
6.7. Case Study
To illustrate SheepDog’s potential for interpretability in real-world scenarios, we conduct a case study based on a fake news sample from the LUN test set. The baseline RoBERTa model correctly identifies the original article as fake, but mistakenly categorizes two style-transformed adversarial articles as real news under LLM-empowered style attacks. In contrast, our SheepDog approach adeptly detects the original article as fake, as well as the two adversarial counterparts. As shown in Figure 3, the article falsely asserts the efficacy of cannabis oil in treating cancer, a claim clearly at odds with established medical knowledge. By leveraging the softmax-converted probabilities derived from the attribution-level prediction scores (refer to Eq. 5), our SheepDog model effectively explains the article’s dubious nature as “false or misleading information”. Combining content-focused veracity attributions with style-agnostic training, SheepDog offers practitioners a valuable tool for reliably discerning and interpreting news veracity.
7. Conclusion and Future Work
In this paper, we address the critical aspect of style-related robustness in fake news detection. Motivated by our empirical finding on the susceptibility of state-of-the-art text-based detectors to LLM-empowered style attacks, we introduce SheepDog, a style-agnostic fake news detector that emphasizes content veracity over style. Leveraging the strengths of task-specific LM backbones and versatile general-purpose LLMs, SheepDog adopts a multi-task learning paradigm, which integrates style-agnostic training and content-focused veracity attributions. Extensive experiments on three benchmark datasets demonstrate SheepDog’s robustness and effectiveness across various style-based adversarial settings, news reframing prompts, and representative LM backbones. Moving forward, SheepDog lays a solid foundation for developing more resilient and adaptable models in the ever-changing online landscape, and demonstrates promising potential to be further extended to multi-modal scenarios.
Appendix A Dataset Description
In Section 6, we conduct extensive experiments on three widely-used real-world benchmark datasets: the FakeNewsNet benchmark (Shu et al., 2020), which consists of the PolitiFact and GossipCop datasets, and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). PolitiFact and LUN contain news articles about political discourse, while GossipCop focuses on celebrity gossip.
In the case of the LUN dataset, which further classifies unreliable news into three sub-categories: satire, hoax, and propaganda, we conduct binary classification between reliable (real) and unreliable (fake) news, and ensure an equal number of unreliable news from each of these fine-grained categories. The detailed statistics for each dataset are shown as follows:
Table 8. Dataset statistics.
Dataset PolitiFact GossipCop LUN
# News Articles 450 7,916 7,500
# Real News 225 3,958 3,750
# Fake News 225 3,958 3,750
Appendix B Details of LLM Prompting
B.1. Prompt Template for Baseline LLM Detectors
In section 6.1.2, we select three representative LLMs as baselines for fake news detection: GPT3.5 (OpenAI, 2022) (version name: gpt-3.5-turbo), InstructGPT (Ouyang et al., 2022) (version name: gpt-3.5-turbo-instruct), and Llama2-13B (Touvron et al., 2023) (version name: Llama-2-13b-chat-hf).
For GPT-3.5 and InstructGPT, we utilize their APIs from OpenAI, and set the temperature to 00 for stable veracity predictions. For Llama2-13B, we utilize the model weights from HuggingFace Transformers (Wolf et al., 2020) version 4.31.0, and set do_sample to False for greedy decoding. For all three LLMs, we adopt the following prompt to conduct zero-shot fake news detection:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Question: Does the following contain real or fake news (or information)? Answer in one word with either ‘Real’ or ‘Fake’, then explain why. [news article]
Answer: [starts with a predicted veracity label]
B.2. Prompt Template for Obtaining SheepDog’s News Reframings
Recall from Section 5.1 that we generate reliable-style reframings and unreliable-style reframings with GPT3.5 for each labeled news article. Among the four prompt templates presented below, we use the first two templates to generate reliable-style reframings, and the other two to generate unreliable-style reframings.
During reframing generation, we set the temperature to $0.7$, and set the maximum number of response tokens to $512$. A detailed reframing example is presented in Table 9.
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in an objective and professional tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a neutral tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in an emotionally triggering tone: [news article]
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a sensational tone: [news article]
B.3. Prompt Template for Obtaining SheepDog’s Veracity Attributions
In Section 5.3, we elicit veracity-related information based on the news content from GPT3.5, specifically by prompting the LLM to assess the labeled articles and their corresponding reframings based on a predefined set of content-focused rationales. We utilize the following prompt template:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Article: [news article]
Question: Which of the following problems does this article have? Lack of credible sources, False or misleading information, Biased opinion, Inconsistencies with reputable sources. If multiple options apply, provide a comma-separated list ordered from most to least related. Answer ”No problems” if none of the options apply.
In the process of prompting GPT3.5 for veracity attributions, we set the temperature to 00 to yield stable attributions, and set the maximum number of response tokens to $300$. We demonstrate an example in Table 10.
Table 9. An example of GPT-3.5 generated reliable-style news reframing.
PROMPT: Rewrite the following article in an objective and professional tone: BREAKING: Obamas Daughter BUSTED By The Cops For Doing This At the Democratic National Convention, Michelle Obama tried to use her daughters as shining examples of just how far America has come. But the fact is that the young Obamas are pushing the limits as any teenager doesand probably even further. According to new reports, Malia Obama was spotted at a party in Marthas Vineyard recently. It was eventually broken up by police because of noise complaints. The first daughter was immediately rushed from the party by Secret Service before the cops arrived. The West Tisbury police declined to confirm if Malia was there, citing national security. But the truth surfaced anyway. This is just one incident in a string of wild behavior on Malias part. Just a month ago the first daughter came under fire when photos surfaced of her smoking marijuana at a concert. Malia is clearly using her gap year before Harvard to party and experiment. Later, Malia and Barack were spotted strolling around a pool together having a chat. Were sure he had a few choice words for her. Clearly, a lack of respect for the law runs deep in this family.
RESPONSE: Malia Obama’s Recent Activities Raise Concerns Recently, reports have emerged regarding the behavior of Malia Obama, daughter of former President Barack Obama and Michelle Obama. At the Democratic National Convention, Michelle Obama showcased her daughters as symbols of progress in America. However, it appears that the young Obamas, like many teenagers, are testing boundaries and possibly going beyond them. According to sources, Malia Obama was seen at a gathering in Marthas Vineyard that was eventually disbanded by law enforcement due to noise complaints. The first daughter was promptly escorted away from the scene by the Secret Service before the arrival of the police. West Tisbury police have neither confirmed nor denied Malia’s presence at the event, citing national security concerns. Nevertheless, the truth regarding her attendance eventually surfaced. This incident is just one in a series of instances where Malia’s behavior has attracted attention. Not long ago, she faced criticism when photos emerged showing her allegedly smoking marijuana at a concert. It appears that Malia is utilizing her gap year before attending Harvard University as a period for socializing and experimentation. Later on, both Malia and Barack Obama were observed engaging in a conversation while walking near a pool. It can be assumed that the former president had some words of advice or guidance for his daughter. It is evident that a lack of regard for the law seems to be a recurring issue within this family. In conclusion, recent events raise concerns about the behavior of Malia Obama. These incidents, including her presence at a party disbanded by the police and alleged marijuana use, suggest a period of exploration and a certain disregard for legal boundaries.
Table 10. An example of content-focused veracity attributions extracted from GPT-3.5.
PROMPT: Article: BREAKING: Obamas Daughter BUSTED By The Cops For Doing This At the Democratic National Convention, Michelle Obama tried to use her daughters as shining examples of just how far America has come. But the fact is that the young Obamas are pushing the limits as any teenager doesand probably even further. According to new reports, Malia Obama was spotted at a party in Marthas Vineyard recently. It was eventually broken up by police because of noise complaints. The first daughter was immediately rushed from the party by Secret Service before the cops arrived. The West Tisbury police declined to confirm if Malia was there, citing national security. But the truth surfaced anyway. This is just one incident in a string of wild behavior on Malias part. Just a month ago the first daughter came under fire when photos surfaced of her smoking marijuana at a concert. Malia is clearly using her gap year before Harvard to party and experiment. Later, Malia and Barack were spotted strolling around a pool together having a chat. Were sure he had a few choice words for her. Clearly, a lack of respect for the law runs deep in this family.
Question: Which of the following problems does this article have? Lack of credible sources, False or misleading information, Biased opinion, Inconsistencies with reputable sources. If multiple options apply, provide a comma-separated list ordered from most to least related. Answer ”No problems” if none of the options apply.
RESPONSE: False or misleading information, Biased opinion.